1. Suppose X is normally distributed with a mean of 5 and a standard deviation of 0.4. Using the standard score formula, we find P (X ≤ X0) = P (Z ≤ 1.3). What is the value of X0?



2. What will be the value of P (X ≥ 235) given that a normal distribution has µ = 241 and d = 2?



3.The average hourly wage of workers at a fast food restaurant is P338.10/hr. Assume the wages are normally distributed with a standard deviation of P23.41. If a worker at this fast food restaurant is selected at random, what is the probability that the worker earns more than P351.10?

1
Expert's answer
2022-06-15T14:01:25-0400

Solution:

1)

"\mu=5;" - mean;

"\sigma=0.4" - standard deviation;

Let's find "X" at "z=1.3;"

"z=\frac{X-\mu}{\sigma};"

"X=z\sigma+\mu=1.3\times0.4+5=5.52;" So, "z=1.3" score 0.4032. It means 40.32% of X values is less than 5.52. And when "X<5.52;" then "z<1.3". So, the result:

"X_0=5.52;"

2)

"\mu=241;" "\sigma=2;" "X_0=235;"

Let"s find "z" :

"z=\frac{X_0-\mu}{\sigma}=\frac{235-241}{2}=-3;" when "z=-3" the score equal 0.0013. It means 0.13% of X values is less than 235. So, "1-0.0013=0.9987;"

"P(X\ge235)=99.87\%."

3)

"\mu=P338.1/h;"

"\sigma=P23.41;"

"X_0=351.1;"

Let's find "z:"

"z=\frac{X_0-\mu}{\sigma}=\frac{351.1-338.1}{23.41}=0.55;" when "z=0.55" the score 0.7088. It means 70.88% of workers' earn less than P351.1. So, "1-0.7088=0.2912;"

"P(X>351.1)=29.12\%;"

Answer:

1) "X_0=5.52;"

2) "P(X\ge235)=99.87\%;"

3) "P(X>351.1)=29.12\%."